# A new class of general linear method with inherent quadratic stability for solving stiff differential systems

### v1

@@ Abstract.
This article proposes a new class of general linear method with $p=q$ and $r=s=p+1$. The construction of the present method is carried out using order conditions and error minimization subject to $A$- stability constraints. The proposed time integration schemes are $A$- and $L$-stable general linear methods (GLMs) equipped with inherent quadratic stability (IQS) criteria. We construct implicit GLMs of orders up to four with $p=q$ and $s=r$ along with the Nordsieck input vector assumption. Further, we test these schemes on three real-world problems: the van der Pol oscillator and two partial differential equations consisting of diffusion (Burgers’ equation and the Gray-Scott model), and numerical results are presented. Computational results confirm that our proposed schemes are competitive with the existing GLMs and can be recognized as an alternative time integration scheme. We demonstrate the order of accuracy and convergence for the proposed schemes through observed order computation and error versus step size plots.
@@
General linear method; Inherent quadratic stability; Nordsieck vector; Stiff differential systems; Burgers’ equation; Gray-Scott model
@@ 1. Introduction
Many important physical and chemical process are described by partial differential equations (PDEs). Examples include transport phenomena, reaction-diffusion systems, and fluid dynamics. These PDEs, upon spatial discretization, result in large systems of ordinary differential equations [1]. However, these differential systems often exhibit complex phenomena that require a suitable time integrator to preserve the solution behavior. The system of time-dependent differential equations in the autonomous form is given by [2]
@@
where, $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with the assumption that (1.1) is well-posed and has a unique solution. In the present work, the partial differential equation governing physical phenomena such as advection-diffusion and reaction-diffusion has been considered. To apply time integration techniques to such a PDE, we need to reduce it to what a time integrator identifies, which is (1.1) using spatial discretization. Choosing an appropriate time integrator becomes crucial [3]. General Linear Methods (GLMs) provide a flexible framework that includes Runge–Kutta (RK) and linear multistep methods (LMM) as special cases [2, 4]. Recent developments in GLMs have produced new classes that are not only accurate but also reliable for such problems [5].
The development of temporal integration schemes based on GLMs has already seen several advancements; GLM and it variant to solve a variety of different problems [6, 7, 8], for more on GLMs, see [9, 10], and the references cited there in. Jackiewicz [9] and Conte et al. [11] built two-step RK and GLM versions with inherent quadratic stability (IQS). Cardone and co-authors [10, 12, 13] studied IQS-based GLMs for non-stiff systems. Later, Braś and his co-authors applied these ideas to stiff-problems [14, 15, 16]. More recently, efficient GLMs with quadratic stability (QS) functions have been constructed to lower the cost of stage computation where s=p1s=p-1 [17]. However, the implicit GLMs with IQS for stiff problems for $s=p+1$ have not been constructed or analyzed.
@@ 2. Construction of implicit GLM with IQS based on Nordsieck input vector
For the initial value problems given by (1.1), the GLM class for one time step $t_{n}=t_{n-1}+h$ is given as follows [2, 9]:
@@
where $\otimes$’ is the Kronecker product and $\mathbf{I}\in\mathbb{R}^{d\times d}$. Here, the components $Y^{[n]}\in\mathbb{R}^{sd\times 1}$ and $y^{[n]}\in\mathbb{R}^{rd\times 1}$ are the internal and external stage approximations
@@
$\mathbf{A}=\begin{bmatrix}\lambda&0&0&\cdots&0\\ a_{21}&\lambda&0&\cdots&0\\ a_{31}&a_{32}&\lambda&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{r1}&a_{r2}&a_{r3}&\cdots&\lambda\end{bmatrix},\quad\mathbf{V}=\begin{bmatrix}1&v_{12}&v_{13}&\cdots&v_{1r}\\ 0&0&v_{23}&\cdots&v_{2r}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&v_{r-1,r}\\ 0&0&0&\cdots&0\end{bmatrix}$.
@@
The internal and external stage approximations, i.e., $Y^{[n]}$ and $y^{[n]}$ respectively, satisfies the following order assumptions [2, 9, 10, 17]
@@
    $y^{[n]}=\sum_{j=0}^{p}\mathbf{l}_{j}h^{j}y^{(j)}(t_{n})+\mathcal{O}(h^{p+1}).$ (2.5)
@@
For the input vector in the Nordsieck form, the vectors $\mathbf{l}_{j}$ become the standard basis of $\mathbb{R}^{r}$, i.e., $\mathbf{l}_{0}=e_{1},\mathbf{l}_{1}=e_{2},\cdots,\mathbf{l}_{p}=e_{r}$.
@@ 2.1. Order conditions
Assume that the GLM has input vector in the Nordsieck form, $y^{[n-1]}=z(t_{n-1})+\mathcal{O}(h^{p+1})$ i.e., $y^{[n-1]}$ is an approximation of the Nordsieck vector of order $p$ with $y_{j}(t_{n-1})=h^{j-1}y^{(j-1)}(t_{n-1})+\mathcal{O}(h^{p+1})$ for $j=1,2,\ldots,r$. The Nordsieck vector is given as follows:
@@ 2.2. Imposition of inherent quadratic stability and minimum error constant
for $\Re(\zeta)<0$ and $|\zeta|$ large, this problem is stiff in nature.
@@
and $\hat{p}(\eta,\omega)$, the stability function is the characteristic polynomial of the matrix $\mathbf{M}(\omega)$.
@@
This matrix equivalence relation for the so-called “inherent quadratic stability (IQS)” condition defined in (2.13) is a sufficient condition for obtaining the stability polynomial in the form given by (2.12). We use the order conditions as defined in Theorem 2.2 and determine the coefficients for the matrix $\mathbf{U}$ and for the r1r-1 columns of $\mathbf{B}$. Now, for determining the rest of the parameters, we impose the algebraic conditions of IQS as defined in (2.13) with the assumption that the matrix $\mathbf{X}$ has the following form [12]
@@
Since, $\tilde{\mathbf{M}}(\omega)\equiv\mathbf{V}$ except for the first two rows we have, $\tilde{\mathbf{M}}_{22}(\omega)=\tilde{\mathbf{V}}_{22}$, and therefore $\tilde{\mathbf{M}}_{22}(\omega)$ is a strictly upper triangular matrix with eigen value 0 of multiplicity r2r-2 and the stability function of matrix $\mathbf{M}(\omega)$ becomes
@@
where, $\tilde{\mathbf{I}}\in\mathbb{R}^{2\times 2}$ and (2.17) gives the required stability function $p(\eta,\omega)$ of the form (2.15).
We search for implicit methods that are $A$-stable as well as $L$-stable, and has possibly a small error constant. Our strategy in this paper is first to employ the $L$-stability criteria given by (2.18) to make the rest of the construction a bit simpler and reduce the number of parameters to compute. For the stability function $p(\eta,\omega)$ given by (2.15) to be $L$-stable, the coefficients $p_{1r}$, and $p_{0r}$ in (2.16) must satisfy the following:
@@
Assuming, $y_{1}^{[n-1]}\approx y(t_{n-1})$, and $y_{j}^{[n-1]}\approx h^{j-1}y^{(j-1)}(t_{n-1})$ for $j=2,3,\cdots,r$ of order $p$ .The error propagation in one step of the GLM scheme is given as follows [10, 12]:
@@ 3. Example Methods
We provide four example methods of the numerical integrators constructed using the order conditions defined in (2.6) and the stability imposition presented in Section 2.2. The proposed classes of implicit GLMs with IQS are $A$- and $L$- stable.
@@ 3.1. GLMQS-1 with $\mathbf{p}=\mathbf{q}=\bf{1}$ and $\mathbf{r}=\mathbf{s}=\bf{2}$
We then solve the minimization problem with constraints obtained by the AA- stability criteria to obtain an example method with an error constant, $E=0.22741$, and the value of the free parameters $v_{12}$ and $\lambda$. The GLM coefficients can then be expressed in the Butcher tableau as follows:
@@
The stability function having two non-zero eigenvalues is given as follows:
@@ 3.2. GLMQS-2 with $\mathbf{p}=\mathbf{q}=\bf{2}$ and $\mathbf{r}=\mathbf{s}=\bf{3}$
Here, the error constant $E=0.0195824$, with the stability function
@@ 3.4. GLMQS-4 with $\mathbf{p}=\mathbf{q}=\bf{4}$ and $\mathbf{r}=\mathbf{s}=\bf{5}$
The error constant for this method is given as $E=2.25574\times 10^{-8}$ and the corresponding stability function is given as follows:
@@ 4. Numerical Experiments
In this section, we provide numerical illustrations for the GLM solvers constructed in Section 3 for three well-known test problems used in the modeling of various real-world phenomena. Firstly, the van der Pol problem, a nonlinear oscillator equation used to model problems in various fields such as electronics, physics, biology, and acoustics. One of its main applications is the simulation of the human heartbeat, particularly in cases where the damping coefficient is large. We also run numerical simulations for Burgers’ equation, a famous PDE that arises in many fields, such as gas dynamics, nonlinear acoustics, fluid mechanics, and traffic flow, among others. Finally, we solve the Gray-Scott model, which is a reaction-diffusion system that shows how chemical processes help move, react, and spread out elements. The Gray Scott model is connected to many areas of physical relevance and may be employed in fields such as biology, epidemiology, medicine, and so on [23]. The effective order of accuracy of the proposed methods is estimated using the formula:
@@
where $h_{1},h_{2}$ are the step sizes. All the order calculations for Tables 1 and 2 are done using (4.1). To assess the accuracy and efficiency of the proposed solvers, we compute error norms from numerical simulations while refining the time grid. The error associated with the step size $h$ is denoted by $e_{h}$. For any general method of order $p$, the global error is expected to behave as follows:

### v2

@@ Abstract.
This article proposes a new class of general linear method (GLM) with $p=q$ and $r=s=p+1$. The construction of the present method is carried out using order conditions and error minimization subject to A- stability constraints. The proposed time integration schemes are A- and L-stable GLMs equipped with inherent quadratic stability (IQS) criteria. We construct implicit GLMs of orders up to four with $p=q$ and $s=r$ along with the Nordsieck input vector assumption. Further, we test these schemes on three real-world problems: the van der Pol oscillator and two partial differential equations (Burgers’ equation and the Gray-Scott model), and numerical results are presented. Computational results confirm that our proposed schemes are competitive with the existing GLMs and can be recognized as an alternative time integration scheme. We demonstrate the order of accuracy and convergence for the proposed schemes through observed order computation and error versus step size plots.
@@
Implicit general linear method; Nordsieck vector; Inherent quadratic stability; Error minimization; Stiff differential systems
@@ 1. Introduction
Many important physical and chemical process are described by partial differential equations (PDEs). Examples include transport phenomena, reaction-diffusion systems, and fluid dynamics. These PDEs, upon spatial discretization, result in large systems of ordinary differential equations (ODEs) [1]. However, these differential systems often exhibit complex phenomena that require a suitable time integrator to preserve the solution behavior. The system of time-dependent differential equations in the autonomous form is given by [2]
@@
where, $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with the assumption that (1.1) is well-posed and has a unique solution. In the present work, the partial differential equation governing physical phenomena such as advection-diffusion and reaction-diffusion has been considered. To apply time integration techniques to such a PDE, we need to reduce it to what a time integrator identifies, which is (1.1) using spatial discretization. But, in problems arising from real-world phenomena, this system is usually stiff, i.e., it requires a range of suitably small step sizes for explicit numerical solvers to obtain stable and convergent solutions, which in turn leads to large computational time and effort. To handle (1.1) when it exhibits stiffness, implicit solvers are typically employed. Therefore, choosing an appropriate time integrator becomes crucial [3]. General Linear Methods (GLMs) provide a flexible framework that includes Runge–Kutta (RK) and linear multistep methods (LMM) as special cases [2, 4]. Recent developments in GLMs have produced new classes that are not only accurate but also reliable for such problems [5].
The development of temporal integration schemes based on GLMs has already seen several advancements; GLM and its variants are used to solve a variety of different problems [6, 7, 8]. For more on GLMs, see [9, 10], and the references cited therein. In recent years, Jackiewicz [9] and Conte et al. [11] built two-step RK and GLM versions with inherent quadratic stability (IQS). Cardone and co-authors [10, 12, 13] studied IQS-based GLMs for non-stiff systems. Later, Braś and his co-authors applied these ideas to stiff-problems [14, 15, 16]. More recently, efficient GLMs with quadratic stability (QS) functions have been constructed to lower the cost of stage computation where s=p1s=p-1 [17]. However, the implicit GLMs with IQS for stiff problems for $s=p+1$ have not been constructed or analyzed.
@@ 2. Construction of implicit GLM with IQS based on Nordsieck input vector
For the initial value problem (IVP) given by (1.1), the GLM class for one time step $t_{n}=t_{n-1}+h$ is given as follows [2, 9]:
@@
where $\otimes$’ is the Kronecker product and $\mathbf{I}\in\mathbb{R}^{d\times d}$. Here, the components $Y^{[n]}\in\mathbb{R}^{sd}$ and $y^{[n]}\in\mathbb{R}^{rd}$ are the internal and external stage approximations
@@
$\mathbf{A}=\begin{bmatrix}\lambda&0&0&\cdots&0\\ a_{21}&\lambda&0&\cdots&0\\ a_{31}&a_{32}&\lambda&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ a_{r1}&a_{r2}&a_{r3}&\cdots&\lambda\end{bmatrix},\quad\mathbf{V}=\begin{bmatrix}1&v_{12}&v_{13}&\cdots&v_{1r}\\ 0&0&v_{23}&\cdots&v_{2r}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&v_{r-1,r}\\ 0&0&0&\cdots&0\end{bmatrix}$. (2.4)
@@
The internal and external stage approximations satisfy the following order assumptions [2, 9, 10, 17].
@@
    $y^{[n]}=\sum_{m=0}^{p}(\mathbf{l}_{m}\otimes\mathbf{I})h^{m}y^{(m)}(t_{n})+\mathcal{O}(h^{p+1}).$ (2.6)
  • $y^{[n-1]}$ is also an approximation of order $p$
    $y^{[n-1]}=\sum_{m=0}^{p}(\mathbf{l}_{m}\otimes\mathbf{I})h^{m}y^{(m)}(t_{n-1})+\mathcal{O}(h^{p+1}).$ (2.7)
@@
For the input vector in the Nordsieck form, the vectors $\mathbf{l}_{m}$ become the standard basis of $\mathbb{R}^{r}$, i.e., $\mathbf{l}_{0}=e_{1},\mathbf{l}_{1}=e_{2},\cdots,\mathbf{l}_{p}=e_{r}$ [17].
@@ 2.1. Order conditions
Assume that the GLM has input vector in the Nordsieck form, $y^{[n-1]}=z(t_{n-1})+\mathcal{O}(h^{p+1})$ i.e., $y^{[n-1]}$ is an approximation of the Nordsieck vector of order $p$ with $y_{j}(t_{n-1})=h^{j-1}y^{(j-1)}(t_{n-1})+\mathcal{O}(h^{p+1})$ for $j=1,2,\ldots,r$. The Nordsieck vector of order p=r1p=r-1 is given as follows:
@@ 2.2. Imposition of inherent quadratic stability and minimum error constant
for $\Re(\zeta)<0$ and $|\zeta|$ is the stiffness parameter.
@@
and define $\hat{p}(\eta,\omega)$, the stability function as the characteristic polynomial of the matrix $\mathbf{M}(\omega)$, i.e.,
@@
This matrix equivalence relation for the so-called “inherent quadratic stability (IQS)” condition defined in (2.15) is a sufficient condition for obtaining the stability polynomial in the form given by (2.14).
We use the order conditions as defined in Theorem 2.2 and determine the coefficients for the matrix $\mathbf{U}$ and for the r1r-1 columns of $\mathbf{B}$. Now, for determining the rest of the parameters, we impose the algebraic conditions of IQS as defined in (2.15) with the assumption that the matrix $\mathbf{X}$ has the following form [12]
@@
Since $\tilde{\mathbf{M}}(\omega)\equiv\mathbf{V}$ except for the first two rows, we have $\tilde{\mathbf{M}}_{22}(\omega)=\tilde{\mathbf{V}}_{22}$, and therefore $\tilde{\mathbf{M}}_{22}(\omega)$ is a strictly upper triangular matrix with eigen values 0 of multiplicity r2r-2 and the stability function of the matrix $\mathbf{M}(\omega)$ becomes
@@
where, $\tilde{\mathbf{I}}\in\mathbb{R}^{2\times 2}$ and (2.19) gives the required stability function $p(\eta,\omega)$ of the form (2.17). ∎
We search for implicit methods that are $A$-stable as well as $L$-stable, and possibly have a small error constant. For the stability function $p(\eta,\omega)$ given by (2.17) to be $L$-stable, the coefficients $p_{1r}$, and $p_{0r}$ in (2.18) must satisfy the following:
@@
Our strategy in this paper is first to employ the $L$-stability criteria given by (2.20) to reduce the number of parameters to compute. Further, we define the local error at time $t_{n}$, which will then be used as an objective in the minimization problem for determining the remaining method coefficients.
Assuming $y_{1}^{[n-1]}=y(t_{n-1})+\mathcal{O}(h^{p+1})$, and $y_{k}^{[n-1]}=h^{k-1}y^{(k-1)}(t_{n-1})+\mathcal{O}(h^{p+1})$ for $k=2,3,\cdots,r$, the error propagation in one step of the GLM scheme is given as follows [10, 12]:
@@
Algorithm 1 Workflow for construction of an Implicit GLM with IQS and $r=s=p+1$
1:Input $p$ (order of method), $s=p+1$ (no. of internal stages), and $r=p+1$ (no. of external stages).
2:Define symbolic matrices: Initialize the coefficient matrices $\mathbf{A}\in\mathbb{R}^{s\times s}$, $\mathbf{U}\in\mathbb{R}^{s\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times s}$, $\mathbf{V}\in\mathbb{R}^{r\times r}$, and abscissae $\mathbf{c}\in\mathbb{R}^{s}$ with $\mathbf{A}$ and $\mathbf{V}$ as given in (2.4).
3:Choose abscissae: Define $\mathbf{c}=[c_{1},\dots,c_{s}]^{T}$ with
$c_{i}=\frac{i-1}{\,s-1\,},\qquad i=1,\dots,s.$
Construct the Vandermonde matrix of the abscissae $\mathbf{c}$
$\mathbf{C}_{r}=\begin{bmatrix}\mathbf{1}&\mathbf{c}&\mathbf{c}^{2}/2!&\dots&\mathbf{c}^{\,r-1}/(r-1)!\end{bmatrix}\in\mathbb{R}^{s\times r}.$
4:Define matrices: Introduce the upper shift matrix $\mathbf{K}_{r}\in\mathbb{R}^{r\times r}$, and
$\mathbf{E}_{r}=\exp(\mathbf{K}_{r})\in\mathbb{R}^{r\times r}.$
5:Solve order conditions: Solve matrix equations
$\mathbf{C}_{r}=\mathbf{A}\,\mathbf{C}_{r}\,\mathbf{K}_{r}+\mathbf{U},\qquad\mathbf{E}_{r}=\mathbf{B}\,\mathbf{C}_{r}\,\mathbf{K}_{r}+\mathbf{V},$
to determine $\mathbf{U}$ and the first (r1)(r-1) columns of $\mathbf{B}$.
6:Applying IQS constraints: Solve $(\mathbf{B}\mathbf{A}-\mathbf{X}\mathbf{B})_{(3:s,:)}=\bf 0$ and $\big(\mathbf{B}\mathbf{U}-(\mathbf{X}\mathbf{V}-\mathbf{V}\mathbf{X})\big)_{(3:r,:)}=\bf 0,$ for all remaining free parameters in $\mathbf{B}$ and $\mathbf{V}$. Here $\mathbf{X}_{(3:s,:)}=\bf 0$ means $\mathbf{X}$ is zero except for first 22 rows.
7:Enforce L-stability: Form the stability function (2.17) and impose the L-stability conditions by setting :
$p_{1r}=0,\qquad p_{0r}=0,$
as given in (2.18).
8:A-Stability via Schur criterion: Apply the Schur stability criteria to obtain the admissible range of the remaining free parameter(s). Denote this admissible set by $\mathcal{F}_{\mathrm{Schur}}$.
9:Minimize the error constant. Solve the minimization problem
$\min_{\text{free parameters}}E_{r}=\left|\frac{1}{(p+1)!}-\frac{\mathbf{b}^{T}\mathbf{c}^{\,p}}{p!}+\mathbf{v}^{T}\beta\right|,\qquad\text{subject to }\mathcal{F}_{\mathrm{Schur}}.$
10:Return the coefficient matrices $\mathbf{A}$, $\mathbf{U}$, $\mathbf{B}$, $\mathbf{V}$, and the abscissa $\mathbf{c}$.
@@ 3. Example Methods
We provide four example methods of the numerical integrators constructed using the order conditions defined in (2.8) and the stability imposition presented in Section 2.2. The proposed classes of implicit GLMs with IQS are AA- and $L$- stable. The workflow of step-by-step procedure for the construction of example methods is demonstrated in Algorithm 1. Following this procedure, we obtain parameter-dependent families of GLMs with IQS, referred to as GLMQS. The coefficient matrices for GLMQS methods of orders up to $p=4$ are provided in the subsequent subsections.
@@ 3.1. GLMQS-1 with $\mathbf{p}=\mathbf{q}=\bf{1}$ and $\mathbf{r}=\mathbf{s}=\bf{2}$
We then solve the minimization problem subject to the constraints of AA- stability criteria to obtain the values of the free parameters $v_{12}$ and $\lambda$. The GLM coefficients can then be expressed in the Butcher tableau as follows:
@@
The error constant for this example method is $E=0.22741$, and the stability function having two non-zero eigenvalues is given as follows:
@@ 3.2. GLMQS-2 with $\mathbf{p}=\mathbf{q}=\bf{2}$ and $\mathbf{r}=\mathbf{s}=\bf{3}$
Here, the error constant, $E=0.0195824$, with the stability function
@@ 3.4. GLMQS-4 with $\mathbf{p}=\mathbf{q}=\bf{4}$ and $\mathbf{r}=\mathbf{s}=\bf{5}$
The error constant for this method is $E=2.25574\times 10^{-8}$ and the corresponding stability function is given as follows:
@@ 4. Numerical Experiments
In this section, we present numerical illustrations for the GLM solvers constructed in Section 3 using three well-known test problems that model various real-world phenomena. First, we consider the van der Pol problem, a nonlinear oscillator equation used in fields such as electronics, physics, biology, and acoustics. One of its primary applications is the simulation of the human heartbeat, particularly when the damping coefficient is large. We then perform numerical simulations for Burgers’ equation, a classical PDE that arises in gas dynamics, nonlinear acoustics, fluid mechanics, traffic flow, and several other areas. Finally, we solve the Gray–Scott model, a reaction–diffusion system that describes how chemical species move, react, and spread. The Gray–Scott model is relevant in many physical and biological contexts and is used in applications including biology, epidemiology, and medicine [23]. The effective order of accuracy of the proposed methods is estimated using the formula:
@@
where $h_{1},h_{2}$ are the step sizes. All the order calculations for Tables 1, 2 and 3 are done using (4.1). To assess the accuracy and efficiency of the proposed solvers, we compute error norms from numerical simulations while refining the time grid. The error associated with the step size $h$ is denoted by $e_{h}$. For any general method of order $p$, the global error is expected to behave as follows: